# Knowledge-Guided Prompt Learning for Request Quality Assurance in Public Code Review

### v1

@@ 6.2. Tag Recommendation Subtask (RQ1)
Our KP-PCR 0.613 0.438 0.247 0.692 0.793 0.871 0.622 0.543 0.373
(1) Compared to baseline methods, KP-PCR enhances the tag recommendation subtask. Our KP-PCR outperforms baselines by 0.1%-29.5% in the tag recommendation task. In terms of handling code snippets, TagDC (Li et al., 2020) and PORFIT (Nie et al., 2020) methods discard code snippets as noise, while Post2Vec (Xu et al., 2021) and PTM4Tag (He et al., 2022) methods encode code snippets separately using the same approach as text segments. The unified modeling of UniPCR adjusts the model architecture through a prompt learning paradigm. Under the similar architecture, KP-PCR achieves better performance by designing a unique training method for code snippets while maintaining the advantage of the unified modeling method, which allows for modeling multiple subtasks.
@@
(3) Introducing external knowledge is helpful for code snippet representation learning. Compared to UniPCR (Chen et al., 2024), it is evident that knowledge-guided prompt learning is more effective in the tag recommendation subtask. As previously mentioned, CDR4Tag (Li et al., 2023b) constructs code snippets based on a matching module, but its shortcomings lie in the complexity and difficult migration of the module, which obscures the information within the code snippets themselves. UniPCR achieves competitive experimental results by narrowing the gap through task reconstruction, while KP-PCR achieves superior performance by introducing prefix vectors and data flow graph representations of code snippets to learn the professional knowledge within the request content.
Overall, KP-PCR shows improvements across all metrics. Compared to the runner-up method UniPCR, it’s evident that knowledge-guided prompting learning is more effective in the tag recommendation subtask. Moreover, CDR4Tag constructs code snippets based on a matching module, but its drawback lies in the complexity of the module, making it difficult to transfer and obscuring the information inherent in the code snippets themselves. While KP-PCR achieves even better performance by introducing prefix vectors and data flow graph representations of code snippets for learning the professional knowledge embedded in the request content.
@@ 6.3. Effects of Knowledge-Guided Propmt Template Designs (RQ2)
This section explores experiments on the template designs for prompt learning, aiming to investigate the impact of using different designs on the model’s performance. Specifically, this section details the prompt learning designs as shown in Table 7.
(1) Knowledge-Guided Soft Prompt. Due to the abundance of domain-specific knowledge in public code reviews, this expertise is underrepresented in traditional language models due to limited coverage. Via generating domain-specific knowledge prefix vectors, external knowledge from Wikipedia is introduced to obtain professional definitions for each label. The professional definitions of the labels are extracted with a length of up to 50 words, and their representation vectors are set as the prefix vectors used in the experiments of this section.
(2) Hard Prompt Construction. Since manually constructed natural language templates are subject to human influence, for better guiding the modeling of the unified modeling method, the hard prompt template is decomposed into two parts: task definition and task description. The task definition guides the model to understand the meaning referred to by the masked word to be predicted in this task, while the task description represents the contextual background significance of the task itself. The task description indicates that the discrete prompt template only contains task-indicative text, while the task definition adds detailed explanations based on the task description. Irrelevant words are replaced with words unrelated to the task, such as “animal,” “you,” “tell,” etc., instead of the prompt language in the template.
(3) Code Prefix Prompt Construction. Code in this study is encouraged as soft prompt to learn representations through a small continuous vector, using fewer parameters to learn the overall meaning. To investigate the impact of continuous vector generation on the unified modeling method proposed in this study, this section designs experiments using irrelevant words and code snippet generation to explore the performance impact of domain-specific knowledge and prefix vector optimization. The experiment with irrelevant words uses words completely unrelated to code reviews and tasks to generate representation vectors.
Table 7. Explanation of Prompt Template Design
Name Knowledge-Guided Task Descriptive Prompt Code Prefix
(a) $\surd$ Task Description Code
(b) $\times$ Task Description Code
(c) $\times$ Irrelevant Words Code
(d) $\times$ Task Description Irrelevant Words
6.3.1. Request Necessity Prediction Subtask
Table 8 displays the experimental results on the necessity prediction subtask.
@@ 6.3.1. Request Necessity Prediction Subtask
(a) 0.830 0.885 0.785
(b) 0.798 0.817 0.775
(c) 0.745 0.771 0.712
(d) 0.775 0.782 0.767
Comparatively, KP-PCR demonstrates improved performance in the necessity prediction subtask of public code review. As shown in Table 8, compared to the suboptimal template (b), if external knowledge guidance is removed, the results will decrease by 1.3%-7.7%. Additionally, the results of template (c) and (d) compared to template (b) show that the accuracy will be reduced if the text descriptive prompt or code prefix is removed. Building upon the unified modeling proposed by KP-PCR, the soft prompt utilizes prefix vector learning to represent code snippets within the overall context, enabling an understanding of the variations in different code snippets under the review necessity criteria. During prediction, leveraging the representations from the learned code snippet prefix vectors enhances the predictive performance of review necessity, resulting in performance improvements for KP-PCR in this subtask.
6.3.2. Tag Recommendation Subtask
Table 9 displays the results of the ablation experiments on the tag recommendation subtask.
@@ 6.3.2. Tag Recommendation Subtask
(a) 0.622 0.543 0.373
(b) 0.619 0.541 0.372
(c) 0.583 0.510 0.351
(d) 0.569 0.507 0.355
(1) The introduction of external knowledge benefits the model in learning the professional knowledge contained in task data. The method of using external knowledge to initialize prefix vectors demonstrates optimal performance. As previously discussed in the analysis and description of public code review and its data in Sec 2, pre-trained language models lack knowledge of specialized terms in code review tasks (such as “c” and “psthewes”), which limits their performance in this domain. External knowledge-guided prefix tuning, by obtaining definitions of technical terms from Wikipedia and using this text as the initialization for prefix tuning, optimizes code snippet representations effectively, leading to excellent performance.
(2) It performs better by using task-related text to construct soft prompt prefix vectors. The use of unrelated words to generate prefix vectors results in the worst model performance and requires more epochs to achieve better performance in this scenario. This indicates that when initializing prefix vectors with unrelated words, the model needs more learning of its representation to adapt to the overall training objective due to a lack of guidance. Experiments initialized with code snippet representations yield suboptimal results, suggesting that initializing with task-related text enhances the model’s learning of the task. Furthermore, this also suggests that better representing code snippets can leverage more features contained in the code snippets to aid in task learning and training.
6.3.3. Discussion
(1) KP-PCR achieves better performance by optimizing the learning of code snippet representations in two quality assurance service subtasks simultaneously. By introducing external knowledge, the model can understand the specific meanings of terms in the request content. By retrieving and matching the relevant definitions from the knowledge base obtained from Wikipedia with the request content, this information is supplemented into the model to help it learn the domain-specific knowledge, resulting in a significant improvement. The experimental results demonstrate its effectiveness.
(2) External knowledge is effective for learning the semantics in the request content. Comparing the direct use of code snippet text to initialize the prefix vector part of the soft prompt, initializing the prefix vector with external knowledge leads to better performance. The use of unrelated words to generate discrete or continuous vectors has a significant impact on the overall experimental results, indicating that a prompt construction method that is more aligned with semantic understanding benefits model performance. Comparing the results of experiments using code snippet text for initialization, it can be observed that introducing external knowledge leads to performance improvement, but the overall performance enhancement is attributed to the adoption of prefix vectors as a learning method adapted to non-natural language text. The effectiveness of the unified modeling, indicating that hign performance can be achieved by selecting appropriate generation methods for initializing hard and soft prompts based on the characteristics of the data.
6.4. Case Study (RQ3)
Currently, with the progress of the GPT series models, the use of large language models to perform tasks such as summarization, translation, and dialogue has attracted widespread attention. Attributed to their powerful reasoning and generation capabilities, numerous prompt learning methods have been proposed to explore the potential of large-scale language models in practical activities, aiming to better guide them in understanding task requirements. Due to its excellent performance on various basic tasks, we design a case study to investigate the performance of our proposed prompt on large-scale language models.
First, we randomly select 20 request text and their corresponding task labels from our experimental dataset. GPT-3.5 and GPT-4.0 are then tasked to select labels based on task definitions. By altering the prompt descriptions, we investigate their performance in quality assurance tasks in public code review.
Table 10. Prompt Design for GPT-3.5 and GPT4.0
The content of the prompt template
Prompt: I need you help me to do a task which is concluding two subtasks, the one is quality prediction, and another is tag recommendation. Quality prediction is to predict the quality tag with a label “good” or “bad” for a text. The tag recommendation is to choose 10 tags in a set for each text (you can only choose tags in this set). I will give you the set for tag recommendation first, then some texts for these task.
Table 10 shows the prompt template used, while Table 11 presents the prediction results of GPT-3.5 and GPT-4.0 for the tag recommendation subtask. The bold black parts in Table 11 indicate results predicted by GPT models that are not in the real label list, while the bold blue parts represent results predicted correctly.
Table 11. The Results of Case Study
Request Text
Basic web server in Python 3. This project is intended to function as a basic web server with a command-line interface for easy use. It’s written in Python 3, with the help of the cmd module for the CLI itself and the thread module (I know, I know) to run the listening/responding process without interfering with the CLI.
I am wondering (with the exception of thread, which I do intend to fix very soon) how well it conforms to best practices and how readable it is. Performance is not a big issue.
import json
version = ””0.3.0””
settings =
serverSocket = socket.socket(socket.AF_INET, socket.SOCK_STREAM)
serverSocket.setsockopt(socket.SOL_SOCKET, socket.SO_REUSEADDR, 1)
def loadSettings():
        global settings
        with open(’.linprconfig’, ’r’) as configFile:
                settings = json.load(configFile)
Ground truth:
¡python¿¡python3x¿¡multithreading¿¡reinventing the wheel¿¡server¿
GPT-3.5:
¡python¿¡web server¿¡CLI¿¡thread module¿¡json¿¡code review¿¡socket programming¿
GPT-4.0:
¡python¿¡server¿¡python3x¿¡cpp¿¡c¿¡coding-style¿¡java¿¡unix¿¡json¿¡formatting¿
KP-PCR: ¡python¿¡beginner¿¡reinventing the wheel¿¡python3x¿¡cpp¿¡c¿¡java¿¡sqlserver¿
¡formatting¿¡server¿
It can be observed that when using natural language in defining task, GPT-3.5 and GPT-4.0 tend to generate ambiguity in the task definitions. Specifically, this manifests as returning results that do not meet the task’s requirements, such as recommending labels not in the real label list or not returning results regarding quality necessity. In contrast, when designing a more logical task statement the predicted results are more aligned with the actual task requirements.
The case study results of the tag recommendation subtask indicate that GPT demonstrates excellent performance on common knowledge labels such as programming languages (Python, Java, etc.), database knowledge (MySQL), computer systems (Linux). It can predict correctly even without pre-training. However, compared to labels containing more specialized information, GPT often fails to make accurate predictions. For example, labels such as “multithreading” and “¡reinventing the wheel” cannot be well predicted even with reference-based learning. This indicates that despite having powerful pre-trained knowledge and rapid learning abilities, large language models need to correctly understand the semantics of specialized knowledge labels.
Overall, the case study with GPT reveals that learning specialized knowledge is important for language models besides general knowledge. In the domain of public code review, there are many technical terms and information hidden within code snippets. Such characteristics make it a challenge for common large language models to handle tasks in the specialized domain of public code review.
@@ 7.2. Public Code Review
Due to the extensive review work involved, conducting code reviews has become one of the biggest challenges facing the industry (Balachandran, 2013). Specifically, code reviews are labor-intensive because reviewers must manually inspect the code and provide written feedback. To alleviate the cost of review requests, considerable attention and research has been devoted to various automation techniques. For instance, Thongtanunam et al. (Thongtanunam et al., 2022) proposed a code transformation service that automatically converts source code into improved versions that have been reviewed and approved. Kudrjavets et al. (Kudrjavets et al., 2022) investigated opportunities for potential speedups in code by analyzing the switch from manual to automatic merging, which they found could enhance code velocity.
@@
Unlike internal code reviews within teams, public code reviews are implemented in the form of software question-and-answer sites. This form, characterized by a large user base and rapid response, also has features such as high knowledge dissemination. However, it also introduces factors that can affect quality and efficiency, primarily related to developers. Therefore, with the popularity of public code review, it has led to task demands similar to code review to ensure the quality of reviews.
@@
As shown in Table 12, previous relevant studies have relied mainly on internal team data or other thematic site data to accomplish intelligent and automated construction of these two tasks. Since the quality and responsiveness of a public code review are typically determined by developers, this process is often carried out manually, making it difficult to achieve convenience and quality assurance. Therefore, with the popularity of this review format, the demand for a shift from manual to intelligent services in public code reviews has also arisen. The process of ensuring review quality in public code reviews can be seen as a review necessity prediction and a tag recommendation task.
Table 12. Popular Software Engineering Thematic Sites
Name Thematic
StackOverFlow Programming Q&A Exchange Sites
WordPress Open-source Blog Design Sites
Code Review Public Code Review Sites
Github Software Project Hosting Sites
Database Administrator Database Management Q&A Sites
@@
Li et al. (Li et al., 2022) proposed that by systematically considering the tasks involved in the entire review process and modeling them, the time and resource costs of the review process can be greatly reduced. Existing research works mostly employ fine-tuning paradigm modifications of model architectures to ensure the model’s performance in terms of accuracy and other relevant metrics for the tasks. This paradigmatic method can accomplish individual tasks with individual models. Table 13 shows the statistical information on model construction for the aforementioned two subtasks. It can be observed that although different modeling methods are used for different tasks, most existing research is based on pre-trained language models. Therefore, considering that existing works in both subtasks have mostly used pre-trained language models for construction in recent years, it is worth considering designing a unified modeling method to simultaneously address both subtasks and reduce cost requirements.
@@ 7.5. Comparison with Our Preliminary Study
In our preliminary study (Chen et al., 2024), a unified model is proposed for the request necessity prediction subtask and the tag recommendation subtask. The study in this paper represents a significant expansion of that study in several respects. Firstly, we introduce an external knowledge base by soft prompt to generate prefix vectors which guide the model to understand the meaning of the code snippet. Secondly, we design the knowledge-guided prompt learning on request quality assurance task to show that the knowledge guide can further improve the request quality of PCR. Thirdly, we set up a prompt template design study to explore the influence of external knowledge. Further, we analyze the time complexity to prove that our KP-PCR can improve task accuracy while maintaining overall efficiency. Finally, we design a case study to investigate the performance of our prompt templates on large-scale language models. In accordance with these additional results, the entire paper has been thoroughly updated to reflect the new holistic study.
8. Conclusions and Future Work
Through investigating and summarizing the relevant tasks of team code review and software question-and-answer sites, and combining them with the process of public code review, this study explores intelligent service methods for the necessity prediction task and tag recommendation task of public code review from the perspective of developers. Addressing the issue of semantic understanding of professional knowledge in the request content, we introduce external knowledge and adopt a processing and training method that conforms to the semantic structure of code snippets to achieve a more comprehensive learning of professional knowledge. We propose KP-PCR to optimize the learning of professional knowledge in public code review. In public code review, request content includes both text and code snippets. The text contains vocabulary related to professional knowledge, while the unique structure of code snippets contains representations of content strongly related to review quality as the main part of the content. By processing code snippets into data flow graphs and guiding the model to learn professional knowledge with knowledge-guided prefix vectors, KP-PCR achieves significant improvements in both subtasks. In the experimental results of the two subtasks, KP-PCR shows an improvement of 8.3% to 28.8% in the necessity prediction subtask and an improvement of 0.1% to 29.5% in the tag recommendation subtask.
This study has used Wikipedia as a source of external knowledge, and the influence of more different knowledge repositories can be explored in the future. In this study, the model was implemented based on the pre-trained language model BERT, which performed well among baseline methods. Further, our prompt learning can be transferred to more large language models in the future.

### v3

@@ 6.2. Tag Recommendation Subtask (RQ1)
Our KP-PCR 0.626 0.447 0.262 0.701 0.803 0.889 0.647 0.561 0.393
(1) Compared to baseline methods, KP-PCR enhances the tag recommendation subtask. Our KP-PCR outperforms baselines by 1.4%-6.9% in the tag recommendation task. In terms of handling code snippets, TagDC (Li et al., 2020) and PORFIT (Nie et al., 2020) methods discard code snippets as noise, while Post2Vec (Xu et al., 2021) and PTM4Tag (He et al., 2022) methods encode code snippets separately using the same approach as text segments. The unified modeling of UniPCR adjusts the model architecture through a prompt learning paradigm. Under the similar architecture, KP-PCR achieves better performance by designing a unique training method for code snippets while maintaining the advantage of the unified modeling method, which allows for modeling multiple subtasks.
@@
(3) Introducing knowledge guidance is helpful for code snippet representation learning. Compared to UniPCR (Chen et al., 2024), it is evident that knowledge-guided prompt learning is more effective in the tag recommendation subtask. As previously mentioned, CDR4Tag (Li et al., 2023b) constructs code snippets based on a matching module, but its shortcomings lie in the complexity and difficult migration of the module, which obscures the information within the code snippets themselves. UniPCR achieves competitive experimental results by narrowing the gap through task reconstruction, while KP-PCR achieves superior performance by introducing prefix vectors and program dependence graph representations of code snippets to learn the professional knowledge within the request content.
Overall, KP-PCR shows improvements across all metrics. Compared to the runner-up method UniPCR, it’s evident that knowledge-guided prompting learning is more effective in the tag recommendation subtask. Moreover, CDR4Tag constructs code snippets based on a matching module, but its drawback lies in the complexity of the module, making it difficult to transfer and obscuring the information inherent in the code snippets themselves. While KP-PCR achieves even better performance by introducing prefix vectors and program dependence graph representations of code snippets for learning the professional knowledge embedded in the request content.
@@ 6.3. Effects of Knowledge-Guided Propmt Template Designs (RQ2)
This section conducts experiments on the template designs for prompt learning, aiming to investigate the impact of using different template designs on the model’s performance. Specifically, we detail the template designs for prompt learning as shown in Table 7.
(1) Knowledge-Guided Soft Prompt. We explore the impact of different types of knowledge guidance on model performance. Specifically, we analyze the effects of using (1) fine-tuned LLMs (DeepSeek, CodeLlama, LLaMA) and (2) general encyclopedic knowledge (Wikipedia) as knowledge guidance in two subtasks (request necessity prediction and tag recommendation). We fine-tune CodeLlama and LLaMA in the same manner as described in Section 4.3.1, using the same dataset and optimizing the fine-tuning parameters to their best values.
(2) Code Prefix Prompt Construction. We design experiments on different code parsing methods, converting code into DFG or PDG, to analyze their impact on the experimental results.
Table 7. Template Designs for Prompt Learning
Name Knowledge-Guided Code Prefix
Our KP-PCR Fine-tuned DeepSeek-7B PDG
(a) Fine-tuned CodeLlama-7B PDG
(b) Fine-tuned Llama3-8B PDG
(c) Fine-tuned DeepSeek-7B DFG
(d) Wikipedia PDG
(e) Wikipedia DFG
(f) $\times$ DFG
@@
Our KP-PCR 0.848 0.894 0.795
(a) 0.845 0.892 0.793
(b) 0.836 0.889 0.790
(c) 0.838 0.890 0.791
(d) 0.835 0.889 0.791
(e) 0.830 0.885 0.785
(f) 0.745 0.771 0.712
@@
Our KP-PCR 0.647 0.561 0.393
(a) 0.646 0.559 0.392
(b) 0.642 0.551 0.385
(c) 0.644 0.555 0.388
(d) 0.635 0.548 0.379
(e) 0.622 0.543 0.373
(f) 0.619 0.541 0.372
Table 8 shows the experimental results on the necessity prediction subtask. Table 9 shows the results of the ablation experiments on the tag recommendation subtask.
(1) Compared to template (a), (b), and (c), using the fine-tuned DeepSeek as an external knowledge source outperforms the fine-tuned CodeLlama, Llama3, or directly using Wikipedia. The fine-tuned DeepSeek improves performance by adjusting model parameters to the needs of specific tasks and providing more precise knowledge related to code review.
(2) The results of template (c) compared to our KP-PCR and the results of template (d) compared to template(e) show that if the code parsing method is changed from DFG to PDG, the accuracy can be improved by 1.5%. PDG and DFG are two commonly used intermediate representation methods, modeling the structural information of programs from different perspectives. PDG models a program as a graph structure composed of statement nodes, with edges representing the composite relationship of control dependence and data dependence. DFG focuses more on depicting the data transfer paths among variables in the program, where nodes typically represent variables or operation units, and edges represent the flow of data. In the PCR tasks we focus on, the model needs to understand the execution logic of code snippets, judge the context where potential defects may occur, and make reasonable inferences about complex dependency relationships. We find that the representation method based on PDG performs more accurately on PCR tasks. This is mainly attributed to the fact that PDG can simultaneously capture the control structure and data transfer of a program, possessing higher structural integrity and semantic expressiveness, and providing more precise context information.
(3) The results of template (f) show that the accuracy reduces when the external knowledge is removed, indicating that external knowledge plays an important auxiliary role in this task.
6.4. LLM Baselines (RQ3)
Recent advancements in large language models (LLMs) have demonstrated remarkable capabilities in code-related tasks. To rigorously evaluate the effectiveness of our proposed knowledge-guided prompt learning framework, we conducted a comparative study by fine-tuning Llama3-8B, CodeLlama-7B, and DeepSeek-7B for the two subtasks separately on our PCR dataset. These models were selected for their proven adaptability to code comprehension tasks.
The fine-tuning process utilized a training set of 61,688 samples and a validation set of 6,855 samples, with task labels aligned to necessity prediction and tag recommendation subtasks. We employed a learning rate of $2\times 10^{-5}$, a batch size of 8, and trained for 10 epochs using mixed-precision training. Both models were optimized with cross-entropy loss and early stopping based on validation performance.
The performance of the fine-tuned models was benchmarked against our proposed KP-PCR framework. Unlike zero-shot or prompt-based evaluations, this approach allows for a fair comparison by leveraging domain adaptation through fine-tuning. As shown in Table 10, our experiments reveal that in the request necessity prediction subtask, LLMs are behind KP-PCR by 22.4%-11.9%. As shown in Table 11, LLMs lag behind KP-PCR by 34.2% and 12.6% in the tag recommendation subtask. These disparities demonstrate that, while LLMs exhibit baseline competence, their reliance on pure textual patterns fails to address the semantic complexity inherent in PCR tasks.
Table 10. Experiment Results of Request Necessity Prediction Subtask
Fine-Tuned LLM Accuracy F1 Necessary F1 Unnecessary
Llama3-8B 0.654 0.583 0.698
CodeLlama-7B 0.692 0.617 0.709
DeepSeek-7B 0.742 0.763 0.721
Our KP-PCR 0.848 0.894 0.795
Table 11. Experiment Results of Tag Recommendation Subtask
Fine-Tuned LLM Precision Recall F1
@3 @5 @10 @3 @5 @10 @3 @5 @10
Llama3-8B 0.411 0.334 0.215 0.321 0.452 0.575 0.346 0.312 0.259
CodeLlama-7B 0.546 0.391 0.207 0.487 0.601 0.762 0.518 0.463 0.352
DeepSeek-7B 0.583 0.457 0.281 0.535 0.634 0.781 0.529 0.508 0.446
Our KP-PCR 0.626 0.447 0.262 0.701 0.803 0.889 0.647 0.561 0.393
@@ 7.2. Code Review
Due to the extensive review work involved, conducting code reviews has become one of the biggest challenges facing the industry (Balachandran, 2013). Specifically, code reviews are labor-intensive because reviewers must manually inspect the code and provide written feedback. To alleviate the cost of review requests, considerable attention and research has been devoted to various automation techniques.
For instance, Thongtanunam et al. (Thongtanunam et al., 2022) proposed a code transformation service that automatically converts source code into improved versions that have been reviewed and approved. Kudrjavets et al. (Kudrjavets et al., 2022) investigated opportunities for potential speedups in code by analyzing the switch from manual to automatic merging, which they found could enhance code velocity.
Tufano et al. (Tufano, 2023) used pre-trained transformers like T5 to automate specific code review tasks such as revising code and generating review comments, aiming to improve code quality and reduce the time developers spend on manual reviews. Hijazi et al. (Hijazi et al., 2023) proposed an approach to evaluate the quality of modern code reviews using biometric data, such as heart rate variability and pupillary response, to assess reviewers’ cognitive load and predict the effectiveness of bug detection, providing feedback on code regions that may require further review. This approach requires gathering the reviewer’s physiological signals such as Heart Rate Variability (HRV) and pupillary responses, which are difficult to retrieve in public code reviews. Hajari et al. (Hajari et al., 2024) proposed a novel code review recommendation system, SofiaWL, which balances expertise, workload distribution, and turnover risk to improve code review efficiency and mitigate knowledge loss in software projects.
These methods are primarily designed around the comments and code changes of individual reviewers, which may limit their effectiveness in the context of community-driven public code reviews.
7.3. Pubilc Code Review
@@
Unlike internal code reviews within teams, public code reviews are implemented in the form of software question-and-answer sites. Such as StackExchange Code Review, a public, community-driven platform, where developers seek feedback from a broad audience. This form, characterized by a large user base and rapid response, also has features such as high knowledge dissemination. However, it also introduces factors that can affect quality and efficiency, primarily related to developers. Public code reviews on websites often involve more complex and dynamic interactions, including discussions, negotiations, and multiple rounds of revisions.
@@
Since the quality and responsiveness of a public code review are typically determined by developers, this process is often carried out manually, making it difficult to achieve convenience and quality assurance. Therefore, with the popularity of this review format, the demand for a shift from manual to intelligent services in public code reviews has also arisen. The process of ensuring review quality in public code reviews can be seen as a review necessity prediction and a tag recommendation task.
@@
Li et al. (Li et al., 2022) proposed that by systematically considering the tasks involved in the entire review process and modeling them, the time and resource costs of the review process can be greatly reduced. Existing research works mostly employ fine-tuning paradigm modifications of model architectures to ensure the model’s performance in terms of accuracy and other relevant metrics for the tasks. This paradigmatic method can accomplish individual tasks with individual models. Table 12 shows the statistical information on model construction for the aforementioned two subtasks. It can be observed that although different modeling methods are used for different tasks, most existing research is based on pre-trained language models.
@@ 7.6. Comparison with Our Preliminary Study
In our preliminary study (Chen et al., 2024), a unified model is proposed for the request necessity prediction subtask and the tag recommendation subtask. The study in this paper represents a significant expansion of that study in several respects. Firstly, we introduce a knowledge guidance base by soft prompt to generate prefix vectors which guide the model to understand the meaning of the code snippet. Secondly, we design the knowledge-guided prompt learning on request quality assurance task to show that the knowledge guide can further improve the request quality of PCR. Thirdly, we set up a prompt template design study to explore the influence of knowledge guidance. Further, we analyze the time complexity to prove that our KP-PCR can improve task accuracy while maintaining overall efficiency. Finally, we design a case study to investigate the performance of our prompt templates on large-scale language models. In accordance with these additional results, the entire paper has been thoroughly updated to reflect the new holistic study.
8. Conclusions
Through investigating and summarizing the relevant tasks of team code review and software question-and-answer sites, and combining them with the process of public code review, this study explores intelligent service methods for the necessity prediction task and tag recommendation task of public code review from the perspective of developers. Addressing the issue of semantic understanding of professional knowledge in the request content, we introduce knowledge guidance and adopt a processing and training method that conforms to the semantic structure of code snippets to achieve a more comprehensive learning of professional knowledge. We propose KP-PCR to optimize the learning of professional knowledge in public code review. In public code review, request content includes both text and code snippets. The text contains vocabulary related to professional knowledge, while the unique structure of code snippets contains representations of content strongly related to review quality as the main part of the content. By processing code snippets into program dependence graphs and guiding the model to learn professional knowledge with knowledge-guided prefix vectors, KP-PCR achieves significant improvements in both subtasks. In the experimental results of the two subtasks, KP-PCR shows an improvement of 5.3% to 27.9% in the necessity prediction subtask and an improvement of 2.4% to 31.4% in the tag recommendation subtask.
Data Availability
The replication package of this study has been made available at https://github.com/WUT-IDEA/KP-PCR.